A rectangle has a length that is 6ft more than the width.Its perimeter is 100 ft. What is the area?
Answer by algebrapro18(249)

A rectangle has a length that is 6ft more than the width.Its perimeter is 100 ft. What is the area?
l = 6+w
l+w = 100
now substituting l into the second equation
l + w = 100
(6+w) + w = 100
6 + 2w = 100
2w = 94
w = 47
now substituting back into the first equation
l = 6 + w
l = 6 + 47
l = 53
so now we know what the length and with are and we can find the Area of a rectangle using the formula A = lw. I will leave this for you to do.